OpenStudy (anonymous):

Find the distance from the point (0,0,0) to the plane x+y+z=1

OpenStudy (anonymous):

suppose that the point is (a,b,c) ....and u have to find its distance from x+y+z-1=0...u have the formula:: \[\left| a+b+c-1 \right|/(\sqrt{a ^{2}+b ^{2}+c ^{2}})\]

OpenStudy (anonymous):

Yes I know this formula but can you show me another way using scalar method?

OpenStudy (anonymous):

wait ill check,...

OpenStudy (anonymous):

according to my text, it's distance is \[scal _{<1,1,1,>}<1,0,0>\]

OpenStudy (anonymous):

ok